# Coarse-Grained Boltzmann Generators

### v1

@@ 5 Related Work
Generative models for equilibrium molecular sampling. Generative models have emerged as powerful tools for sampling Boltzmann distributions (Olsson, 2026; Klein et al., 2023; Rotskoff, 2024; Aranganathan et al., 2025; Janson and Feig, 2025). In particular, Boltzmann Generators (Noé et al., 2019) represent an influential class of such models. Subsequent work has improved BGs in various ways (von Klitzing et al., 2025; Schebek and Rogal, 2025), including incorporating inductive biases (Köhler et al., 2020) and enhancing transferability across chemical and thermodynamic conditions (Klein and Noé, 2024; Dibak et al., 2022; Moqvist et al., 2025; Invernizzi et al., 2022). Despite efforts to employ scalable architectures (Tan et al., 2025a; Zhai et al., 2024) or reduce the cost of likelihood evaluation (Gloy and Olsson, 2025; Rehman et al., 2025; Peng and Gao, 2025), atomistic BGs remain limited in scalability. CG Boltzmann Emulators (Lewis et al., 2025; Zheng et al., 2024) bypass the reweighting step, enabling larger-scale applications at the cost of asymptotic exactness. To bridge the resolution gap, backmapping (Chennakesavalu et al., 2023; Hummerich et al., 2025; Wang et al., 2022) or other reconstruction strategies (Tamagnone et al., 2024; Schopmans and Friederich, 2024; Stupp and Koutsourelakis, 2025) are often employed to recover atomistic detail. Another active line of research considers neural samplers for sampling unnormalized densities (Akhound-Sadegh et al., 2024; Midgley et al., 2023; He et al., 2025; Potaptchik et al., 2025; Liu et al., 2025), which also have applications in molecular systems (Nam et al., 2025; Havens et al., 2025).
@@
Coarse-grained machine learning potentials. CG potentials have historically been parameterized using either top-down or bottom-up approaches. In top-down methods, model parameters are adjusted to reproduce macroscopic observables, such as experimental measurements (Marrink et al., 2007; Thaler and Zavadlav, 2021; Fuchs and Zavadlav, 2025). In contrast, bottom-up approaches aim to construct CG models that reproduce the marginal equilibrium distribution or free energy landscape of the underlying atomistic system (Jin et al., 2022; Noid, 2013). Prominent bottom-up methods include variational force matching (Noid et al., 2008) and relative entropy minimization (Shell, 2008). Recent progress in machine learning has opened new opportunities for learning CG MLPs (Plainer et al., 2025; Kohler et al., 2023; Arts et al., 2023; Wang et al., 2019; Zhang et al., 2018). The development of CG MLPs can be viewed as an extension of broader effort to construct accurate machine learning interatomic potentials from first-principles calculations. (Batatia et al., 2022; Wood et al., 2025; Batzner et al., 2022; Unke et al., 2021; Kabylda et al., 2025). Despite these successes, learning transferable and efficient CG MLPs remains an open challenge (Charron et al., 2025; Mirarchi et al., 2024; Majewski et al., 2023; Durumeric et al., 2023).
@@ 6 Conclusion
This work introduces CG-BGs, a scalable framework for equilibrium sampling of coarse-grained molecular systems. By targeting the marginal equilibrium distribution defined by the PMF, CG-BGs reduce the effective dimensionality of the sampling problem while preserving thermodynamic consistency through importance reweighting. The underlying PMF can be learned from rapidly converged data using enhanced sampling force matching, providing a correction mechanism for existing CG Boltzmann Emulators. Even at high levels of coarse-graining, CG-BGs capture solvent-mediated and many-body effects, and enable one-shot, simulation-free evaluation of CG MLPs.
Limitations. The current approach relies on predefined collective variables for coarse-graining and enhanced sampling, which may be nontrivial to identify for complex systems. Recent advances in collective variable discovery (Zhang et al., 2024; Ribeiro et al., 2018; Chen and Ferguson, 2018; Herringer et al., 2023; Mehdi et al., 2024) and uncertainty quantification (Zaverkin et al., 2024; Musil et al., 2019) provide promising avenues to address these challenges.
Future work. Extending CG-BGs to larger, more complex systems is a natural next step, leveraging demonstrated transferability of flow-based generative models and MLPs. The simulation-free evaluation paradigm introduced here also opens new opportunities, including integrating CG-BGs with experimental data by incorporating it into the learned PMF, and accelerating differentiable training strategies (Thaler and Zavadlav, 2021; Han and Yu, 2025) by replacing iterative simulations with high-quality generated samples.
Broader Impact
@@ B.2 Alanine Dipeptide
Force Fields. For the explicit solvent dataset, alanine dipeptide is parameterized using the AMBER99SB-ILDN force field (Lindorff-Larsen et al., 2010) and solvated in a cubic box of TIP3P water molecules. All explicit solvent simulations are carried out with GROMACS (Van Der Spoel et al., 2005). For the implicit solvent dataset, the same force field is used together with the generalized Born (OBC1/OBC2) implicit solvent model, and simulations are performed using OpenMM (Eastman et al., 2023).
@@
Well-Tempered Metadynamics. We perform well-tempered metadynamics (WT-MetaD) (Barducci et al., 2008) simulations of alanine dipeptide in explicit solvent using GROMACS coupled with PLUMED (Bonomi et al., 2009). The backbone dihedral angles $\phi$ (C–N–Cα–C) and $\psi$ (N–Cα–C–N) are chosen as collective variables. Gaussian hills with height $1.2\,\text{kJ/mol}$ and width $0.35\,\text{rad}$ are deposited every 500 integration steps ($1\,\text{ps}$). Datasets are generated with bias factors $\gamma=1.5$ and $\gamma=9$, where $\gamma=1.5$ is used to train flow model for biased proposal generation and $\gamma=9$ is used for enhanced sampling force matching. Each simulation is run for $2\times 10^{7}$ steps ($10\,\text{ns}$) with a time step of $0.5\,\text{fs}$ and no bond constraints. Positions and forces are recorded every 40 steps ($0.02\,\text{ps}$), yielding $5\times 10^{5}$ saved configurations per dataset. To ensure unbiased force labels, all forces are recomputed by rerunning the saved trajectories in GROMACS without the metadynamics bias using the mdrun -rerun functionality. This guarantees that each configuration is associated with forces from the underlying unbiased potential.
@@
Simulation Details. All simulations are performed in the NVT ensemble at a temperature of 300 K. For explicit solvent systems, alanine dipeptide is solvated in a cubic box of TIP3P water with a side length of 3.7 nm. After energy minimization, production dynamics are carried out using a velocity-rescale thermostat (time constant 0.1 ps). Long-range electrostatics are treated using the particle mesh Ewald method, and van der Waals interactions are truncated at 1.0 nm. We summarize the dataset configurations used in this work in Tab. 3.
Table 3: Overview of simulation details for alanine dipeptide datasets.
Solvent Dataset Method Length Timestep Constraints
Explicit (TIP3P) Unbiased MD None 500 ns 0.5 fs None
Explicit (TIP3P) Biased MD (for CNF) WT-MetaD ($\gamma=1.5$) 10 ns 0.5 fs None
Explicit (TIP3P) Biased MD (for PMF) WT-MetaD ($\gamma=9$) 10 ns 0.5 fs None
Implicit (OBC1) Unbiased MD None 500 ns 2 fs H-bonds
Implicit (OBC2) Unbiased MD None 500 ns 2 fs H-bonds
@@ C.1 Architecture
Alanine Dipeptide. For the alanine dipeptide, we use an adapted Graph Transformer architecture (Plainer et al., 2025; Arts et al., 2023; Shi et al., 2020). Given bead positions $\mathbf{x}_{i}$ and bead features $\mathbf{h}_{i}$, edge attributes are constructed as
@@
where $t$ denotes the flow time. The bead features and time are embedded into 16- and 4-dimensional vectors, respectively, concatenated with positions, and projected to 128-dimensional node embeddings via a linear layer. Edge attributes are also embedded into 128 dimensions. The model consists of three Graph Transformer layers with 8 attention heads and a head dimension of 64. To enforce rotational equivariance, random global rotations are applied to molecular configurations during training as data augmentation (Abramson et al., 2024).
@@ C.2 Training Configuration
For alanine dipeptide, models are trained using AdamW with weight decay $10^{-5}$ and a cosine learning-rate schedule ranging from $3\times 10^{-4}$ to $1\times 10^{-5}$, using a batch size of 256 for 5000 epochs. For MB potential, we use the same configuration except training for 2000 epochs. We find no consistent benefit from exponential moving average (EMA) of parameters and therefore do not use it in our experiments.
@@ C.3 Inference
As Hutchinson’s trace estimator introduces bias for BGs (Tan et al., 2025a; Peng and Gao, 2025), we compute the divegence exactly using automatic differentiation. For inference, we use the Dormand–Prince 5(4) method (dopri5) with absolute and relative tolerances set to $10^{-4}$. Inference is performed with a batch size of 1000.
@@ C.4 Computational Cost
The reported inference times (Tab. 4) correspond to generating $10^{4}$ samples. All other training and inference parameters are provided in §C.2 and §C.3. A batch size of 1000 is used for all experiments, except for the all-atom mapping, which uses a batch size of 512 due to GPU memory constraints.
Table 4: Training and inference time for CFM across settings. Inference times correspond to $10^{4}$ generated samples. Results marked with are from (Rehman et al., 2025).
Core Beta Heavy Atom All Atom ECNF++ SBG DiT CNF FALCON
Training 0.83h 1.39h 3.93h
Inference 0.84min 3.46min 13.92min
Total 0.84h 1.45h 4.16h 12.52h 16.83h 9.56h 7.65h
@@ D.1 Architecture
Alanine Dipeptide. For alanine dipeptide, the CG potential $U_{\eta}(\mathbf{R})$ is parameterized using the MACE architecture (Batatia et al., 2022), an equivariant message-passing graph neural network. Each CG bead is represented as a node in a geometric graph, with edges connecting neighbors within a cutoff radius and encoding relative position vectors. The model uses hidden irreducible representations of $32\times 0e+32\times 1o$, processed through two interaction layers with correlation order 3 and an angular momentum expansion truncated at $\ell_{\rm max}=3$. Node features are decoded by a readout layer ($16\times 0e$) into a scalar energy prediction ($1\times 0e$). Periodic displacement functions are applied during graph construction to handle boundary conditions correctly.
@@ D.2 Training Configuration
For MB, training uses the Adam optimizer (via Optax) with a constant learning rate of $10^{-4}$ and batch size 128 for 500 epochs. For alanine dipeptide, training uses Adam with exponential learning rate decay ($\eta_{0}=10^{-3}$, decay rate 0.9), batch size 256, and 250 epochs. Training and validation splits are with a 90/10 ratio. Gradients are clipped to a global norm of 1.0. The final parameters for energy evaluation correspond to the minimum validation loss.
@@ D.3 Computational Cost
Table 5: Training and inference time for MACE model. Inference times correspond to $10^{4}$ samples evaluated.
@@ Appendix G Additional Results
G.1 Ramachandran Plots
We show Ramachandran plots for alanine dipeptide to illustrate the effect of importance reweighting. Fig. 6 shows the reference distributions from MD simulations in explicit and implicit solvent. Fig. 6b corresponds to a short 10 ns unbiased trajectory, which fails to sample the correct $\phi$ mode, highlighting the need for enhanced sampling to achieve sufficient configurational coverage.
We compare the MD reference with CG-BG flow proposals before and after reweighting for both mappings (Fig. 7, Fig. 8, Fig. 10, and Fig. 11). Unreweighted proposals exhibit noisy sampling and populate low-probability regions of the free energy landscape. These configurations receive low importance weights and contribute negligibly after reweighting, yielding distributions consistent with the reference.
We further examine reweighted samples obtained using different PMFs. The PMF trained on long unbiased data ($\text{PMF}_{U}$) underestimates the correct metastable basin of the $\phi$ angle (Fig. 9), reflecting the limitations of learning CG PMFs from slowly converging unbiased trajectories.
Figure 6: Ramachandran plots for MD simulations. (a) 500 ns unbiased MD simulation (explicit solvent). (b) 10 ns unbiased MD simulation (explicit solvent) . (c) 500 ns unbiased MD simulation (implicit solvent) (OBC2).
Figure 7: Ramachandran plots (Heavy Atom). (a) MD reference. (b) Flow proposal (flow trained on unbiased dataset). (c) Reweighted result ($\text{PMF}_{B}$).
Figure 8: Ramachandran plots (Heavy Atom). (a) MD reference. (b) Flow proposal (flow trained on WT-MetaD dataset). (c) Reweighted result ($\text{PMF}_{B}$).
Figure 9: Ramachandran plots (Heavy Atom). (a) MD reference. (b) Flow proposal (flow trained on unbiased dataset). (c) Reweighted result ($\text{PMF}_{U}$).
Figure 10: Ramachandran plots (Core Beta). (a) MD reference. (b) Flow proposal (flow trained on unbiased dataset). (c) Reweighted result ($\text{PMF}_{B}$).
Figure 11: Ramachandran plots (Core Beta). (a) MD reference. (b) Flow proposal (flow trained on WT-MetaD dataset). (c) Reweighted result ($\text{PMF}_{B}$).
G.2 Ablation of Weight Clipping

### v2

@@ 5 Related Work
Generative models for equilibrium molecular sampling. Generative models have emerged as powerful tools for sampling Boltzmann distributions (Olsson, 2026; Klein et al., 2023; Rotskoff, 2024; Aranganathan et al., 2025; Janson and Feig, 2025; Xie et al., 2026). Subsequent work has improved BGs in various ways (von Klitzing et al., 2025; Schebek and Rogal, 2025; OuYang et al., 2026), including the incorporation of inductive biases (Köhler et al., 2020), improved transferability across chemical and thermodynamic conditions (Klein and Noé, 2024; Dibak et al., 2022; Moqvist et al., 2025; Invernizzi et al., 2022), and more scalable architectures and likelihood estimators (Tan et al., 2025a; Zhai et al., 2024; Gloy and Olsson, 2025; Rehman et al., 2025; Peng and Gao, 2025). Several approaches address the resolution gap between coarse-grained and atomistic representations through backmapping (Chennakesavalu et al., 2023; Hummerich et al., 2025; Wang et al., 2022) or other reconstruction strategies (Schopmans and Friederich, 2024; Stupp and Koutsourelakis, 2025). Closely related to our setting, Tamagnone et al. (2024) combine a normalizing flow over collective variables with nonequilibrium dynamics to evolve the remaining degrees of freedom. Kohler et al. (2023) instead use a normalizing flow to model the coarse-grained distribution and generate configurations and forces for training coarse-grained MLPs via force matching. Another active line of research considers neural samplers for sampling unnormalized densities (Akhound-Sadegh et al., 2024; Midgley et al., 2023; He et al., 2025; Potaptchik et al., 2025; Liu et al., 2025), which also have applications in molecular systems (Nam et al., 2025; Havens et al., 2025; Blessing et al., 2026).
@@
Coarse-grained machine learning potentials. The development of CG MLPs can be viewed as an extension of broader effort to construct accurate machine learning interatomic potentials from first-principles calculations. (Batzner et al., 2022; Batatia et al., 2022; Unke et al., 2021). Beyond bottom-up approaches (Jin et al., 2022; Noid, 2013), CG potentials can also be parameterized using top-down methods that reproduce macroscopic observables or experimental measurements (Marrink et al., 2007; Thaler and Zavadlav, 2021; Fuchs and Zavadlav, 2025). Recent work has explored different machine learning approaches for learning CG MLPs (Zhang et al., 2018; Wang et al., 2019; Kohler et al., 2023; Arts et al., 2023; Plainer et al., 2025). Despite these advances, learning transferable and computationally efficient CG MLPs remains challenging (Charron et al., 2025; Mirarchi et al., 2024; Majewski et al., 2023; Durumeric et al., 2023). A limitation of standard force matching objectives is their strong reliance on large amounts of converged simulation data. ESFM (Chen et al., 2026) addresses this by learning from rapidly converged biased simulations. Alternative strategies include constrained MD approaches for approximating mean forces prior to training (Ciccotti et al., 2005; Park et al., 2026; Fan et al., 2026).
@@ 6 Conclusion
This work introduces CG-BGs, a scalable framework for equilibrium sampling of coarse-grained molecular systems. By targeting the marginal equilibrium distribution defined by the PMF, CG-BGs reduce the effective dimensionality of the sampling problem while retaining asymptotic correctness through importance reweighting. The underlying PMF can be learned from rapidly converged data using enhanced sampling force matching, providing a correction mechanism for existing CG Boltzmann Emulators. Even at high levels of coarse-graining, CG-BGs capture solvent-mediated and many-body effects, and enable one-shot, simulation-free evaluation of CG MLPs.
Limitations. The current approach uses predefined collective variables for coarse-graining and enhanced sampling, which may be nontrivial to identify for complex systems. Recent advances in collective variable discovery (Zhang et al., 2024; Ribeiro et al., 2018; Chen and Ferguson, 2018; Herringer et al., 2023; Mehdi et al., 2024) and uncertainty quantification (Zaverkin et al., 2024; Musil et al., 2019) provide promising avenues to address these challenges.
Future work. Extending CG-BGs to larger, more complex systems is a natural next step, leveraging demonstrated transferability of generative models (Didi et al., 2026; Antoniadis et al., 2026) and MLPs (Wood et al., 2025; Kabylda et al., 2025). More broadly, advances in exact-likelihood generative modeling, including autoregressive architectures (Rehman et al., 2026; Yu et al., 2026), as well as improvements in atomistic Boltzmann Generators, can be readily transferred to the CG setting. The simulation-free evaluation method introduced here also opens the possibility of elucidating the design space of coarse-grained machine learning potentials. It could enable systematic assessment of how architectural choices, parameterizations, and training objectives influence PMF accuracy. In contrast to conventional pipelines that rely on validation loss or repeated MD simulations, this approach could provide direct assessment of model quality via observable level comparisons. Finally, while CG-BGs are trained from simulation data in this work, one could also explore energy-based training or more general neural sampler formulations using the learned PMFs as unnormalized targets.
@@ B.2 Alanine Peptides
Force Fields. For the explicit solvent dataset, alanine peptide systems are parameterized using the AMBER99SB-ILDN force field (Lindorff-Larsen et al., 2010) and solvated in a cubic box of TIP3P water molecules. Explicit solvent simulations of alanine dipeptide are carried out with GROMACS (Van Der Spoel et al., 2005), while all remaining peptide systems are simulated using OpenMM (Eastman et al., 2023). For implicit solvent, the same force field is used together with the generalized Born (OBC1/OBC2) model, and all simulations are performed using OpenMM.
@@
Well-Tempered Metadynamics. We perform well-tempered metadynamics (WT-MetaD) (Barducci et al., 2008) simulations of alanine peptides in explicit solvent using GROMACS coupled with PLUMED (Bonomi et al., 2009). The backbone dihedral angles $\phi$ (C–N–Cα–C) and $\psi$ (N–Cα–C–N) are chosen as collective variables. Gaussian hills with height $1.2\,\text{kJ/mol}$ and width $0.35\,\text{rad}$ are deposited every 500 integration steps. Datasets are generated with bias factors $\gamma=1.5$ and $\gamma=9$, where $\gamma=1.5$ is used to train flow model for biased dipeptide proposal generation and $\gamma=9$ is used for enhanced sampling force matching. Positions and forces are recorded. To ensure unbiased force labels, all forces are recomputed by rerunning the saved trajectories in GROMACS without the metadynamics bias using the mdrun -rerun functionality. This guarantees that each configuration is associated with forces from the underlying unbiased potential.
For WT-MetaD, the backbone dihedral pair $(\phi,\psi)$ is used as collective variables for the dipeptide system. For the tripeptide and hexapeptide systems, biasing is applied only to the specific dihedral pair of interest, which is the second pair (of three) for alanine tripeptide and the third pair (of six) for alanine hexapeptide, counting from the N-methyl terminus. As a result, convergence is primarily enforced along the targeted collective variables, which we find sufficient for accurate estimation of the PMF along the corresponding degrees of freedom in this work. More general biasing schemes or longer simulations may further improve sampling of the remaining degrees of freedom, and thereby improve the accuracy of the global PMF.
@@
Simulation Details. All simulations are performed in the NVT ensemble at a temperature of 300 K, with a time step of $0.5\,\text{fs}$ and no bond constraints. After energy minimization, production dynamics are carried out using a velocity-rescale thermostat (time constant 0.1 ps). Long-range electrostatics are treated using the particle mesh Ewald method, and van der Waals interactions are truncated at 1.0 nm. We summarize the dataset configurations used in this work in Tab. 4.
Table 4: Overview of simulation details for alanine peptide datasets.
System Solvent Dataset Method Length
Dipeptide Explicit (TIP3P) Unbiased MD None 500 ns
Explicit (TIP3P) Biased MD (CNF) WT-MetaD ($\gamma=1.5$) 10 ns
Explicit (TIP3P) Biased MD (PMF) WT-MetaD ($\gamma=9$) 10 ns
Implicit (OBC1) Unbiased MD None 500 ns
Implicit (OBC2) Unbiased MD None 500 ns
Tripeptide Explicit (TIP3P) Unbiased MD None 1000 ns
Explicit (TIP3P) Biased MD (PMF) WT-MetaD ($\gamma=9$) 50 ns
Implicit (OBC2) Unbiased MD None 1500 ns
Hexapeptide Explicit (TIP3P) Unbiased MD None 1500 ns
Explicit (TIP3P) Biased MD (PMF) WT-MetaD ($\gamma=9$) 100 ns
Implicit (OBC2) Unbiased MD None 1500 ns
@@ C.1 Architecture
Alanine Peptides. For peptides, we use an adapted Graph Transformer architecture (Plainer et al., 2025; Arts et al., 2023; Shi et al., 2020). Given bead positions $\mathbf{x}_{i}$ and bead features $\mathbf{h}_{i}$, edge attributes are constructed as
@@
where $t$ denotes the flow time. The bead features and time are embedded into 16- and 4-dimensional vectors, respectively, concatenated with positions, and projected to 128-dimensional node embeddings via a linear layer. Edge attributes are also embedded into 128 dimensions. The model consists of three Graph Transformer layers with 8 attention heads and a head dimension of 64. To enforce rotational equivariance, random global rotations are applied to molecular configurations during training as data augmentation (Abramson et al., 2024). Additionally, translational equivariance is guaranteed by moving the center of mass to the origin and adding noise to the center of mass to lift the data dimensionality back. (Tan et al., 2025a).
@@ C.2 Training Configuration
All models are trained using AdamW with weight decay $10^{-5}$ and a cosine learning-rate schedule decreasing from $3\times 10^{-4}$ to $1\times 10^{-5}$. For the Müller–Brown potential, training is performed for 1000 epochs with batch size 256 using 20k training samples. For alanine dipeptide, models are trained for 5000 epochs with batch size 1024 using 50k samples. For alanine tripeptide and hexapeptide, models are trained for 10000 epochs with batch size 1024 using 200k samples. We find no consistent benefit from exponential moving average (EMA) of model parameters and therefore do not employ it in our experiments.
@@ C.3 Inference
As Hutchinson’s trace estimator introduces bias for BGs (Tan et al., 2025a; Peng and Gao, 2025), we compute the divergence exactly using automatic differentiation. For inference, we use the Dormand–Prince 5(4) method (dopri5) with absolute and relative tolerances set to $10^{-5}$. Inference is performed with a batch size of 500.
@@ C.4 Computational Cost
The reported inference times (Tab. 5) correspond to generating $10^{4}$ samples. All other training and inference parameters are provided in §C.2 and §C.3.
Table 5: Training and inference time for CFM across different systems and settings. Inference times correspond to $10^{4}$ generated samples. Results marked with are taken from (Rehman et al., 2025) and are included for reference. Reported runtimes should be interpreted in light of differences in hardware, implementation details, and training and inference batch sizes. Note also that peptide nomenclature differs between the two works due to different conventions for counting terminal capping groups: our alanine tripeptide and hexapeptide correspond to the alanine tetrapeptide and heptapeptide systems reported in (Rehman et al., 2025), respectively.
Model Alanine Dipeptide Alanine Tripeptide
Training Inference Total Training Inference Total
Core Beta 0.45h 0.95min 0.47h 7.32h 6.47min 7.43h
Heavy Atom 0.80h 3.78min 0.86h 17.63h 14.22min 17.87h
All Atom 2.55h 14.91min 2.80h
ECNF++ 12.52h 32.17h
SBG 16.83h 41.67h
DiT CNF 9.56h 24.10h
FALCON 7.65h 25.76h
Model Alanine Hexapeptide
Training Inference Total
Core Beta 23.01h 20.07min 23.34h
@@ D.1 Architecture
Alanine Peptides. For peptides, the CG potential $U_{\eta}(\mathbf{R})$ is parameterized using the MACE architecture (Batatia et al., 2022), an equivariant message-passing graph neural network. Each CG bead is represented as a node in a geometric graph, with edges connecting neighbors within a cutoff radius and encoding relative position vectors. The model uses hidden irreducible representations of $32\times 0e+32\times 1o$, processed through two interaction layers with correlation order 3 and an angular momentum expansion truncated at $\ell_{\rm max}=3$. Node features are decoded by a readout layer ($16\times 0e$) into a scalar energy prediction ($1\times 0e$). Periodic displacement functions are applied during graph construction to handle boundary conditions correctly.
@@ D.2 Training Configuration
For MB, training uses the Adam optimizer (via Optax) with a constant learning rate of $10^{-4}$ and batch size 128 for 500 epochs. For peptides, training uses Adam with exponential learning rate decay ($\eta_{0}=10^{-3}$, decay rate 0.01), batch size 256, and 300 epochs. We use 500k training samples for alanine dipeptide and 1000k training samples for alanine tripeptide and hexapeptide across all CG mappings. Training and validation splits are with a 90/10 ratio. Gradients are clipped to a global norm of 1.0. Validation losses in CG force matching are often noisy and do not consistently correlate with potential quality. We use the final training checkpoint for inference in all experiments.
@@ D.3 Computational Cost
Table 6: Training and inference time for MACE model on alanine dipeptide. Inference times correspond to $10^{4}$ samples evaluated.
@@ Appendix G Additional Results
G.1 Müller-Brown Potential
Figure 6: CG-BGs on the MB potential. (a) Two-dimensional unbiased MB potential energy surface (functional form in §B). (b) Marginal probability density along the $x$ coordinate. (c) Free energy profiles before and after reweighting for CG-BGs, where flow is trained on unbiased data, compared with the exact solution and MD reference. (d-f) Same as (a-c), but for flow trained on biased data.
G.2 Alanine Tripeptide (Core Beta mapping)
Figure 7: CG-BGs on alanine tripeptide (Core Beta). (a) Core Beta mapping. (b) Potential energy distribution under the learned PMF. (c) $\phi$ dihedral free energy profile. (d) MD reference Ramachandran plot. (e) Reweighted Ramachandran distribution obtained from CG-BGs.
G.3 Ramachandran Plots
We show Ramachandran plots for peptides to illustrate the effect of importance reweighting (Fig. 8 and Fig. 9), using MD reference distributions from explicit and implicit solvent simulations. The raw flow proposals exhibit noisy sampling and place probability mass in low-probability regions of the free energy landscape. These configurations receive low importance weights and therefore contribute negligibly after reweighting, yielding distributions that closely match the explicit solvent reference MD ensembles.
Figure 8: Ramachandran plots for alanine dipeptide. (a,g) Reference MD (explicit solvent). (b) Implicit solvent MD (OBC1). (h) Implicit solvent MD (OBC2). (c) Core Beta (unbiased) proposal. (d) Core Beta (WT-MetaD) proposal. (e) Heavy Atom (unbiased) proposal. (f) Heavy Atom (WT-MetaD) proposal. (i–l) Reweighted distributions corresponding to (c–f).
Figure 9: Ramachandran plots for alanine tripeptides and hexapeptides. (a,b) Reference MD for tripeptide (explicit and implicit solvent). (c,d) Flow proposals for tripeptide (Core Beta and Heavy Atom). (e) Flow proposal for hexapeptide (Core Beta). (f,g) Reference MD for hexapeptide (explicit and implicit solvent). (h–j) Reweighted distributions corresponding to (c–e).
G.4 Ablation of Weight Clipping